Demand is described by the equation: Qd = 200-5P, and supply: Qs = 35P-200. The government introduced a per-product tax on the producer in the amount of t per unit of goods. As a result, the equilibrium quantity (q) decreased by 17,5%. Find the amount of tax (t)


1
Expert's answer
2021-11-26T12:24:46-0500

At equilibrium Qd=Qs

"200-5P = 35P -200 \\

400 = 40P \\

P = 10 \\

Q = 200 -5 \times 10 = 150"

New "Q^* = (150 - 150 \times 0.175) = 150 -26.25 = 123.75"

"123.75 = 200 -5P^* \\

5P^* = 76.25 \\

P^* = 15.25"

New price = 15.25

The amount of tax

ΔP = 15.25-10=5.25